A body is dropped from a balloon moving with velocity 4m/s when the balloon is at a height of 120.5 m from the ground find the height of the body after 5 seconds from the ground

Grade:11

3 Answers

Vikas TU
7 years ago
body relative vel. = 0 – 4j (when it is dropped).
distance travelled by it,
x = 4*5 - 0.5g*25
 = 20 - 125
 = -105 meter
Height of the body from ground = 120.5 – 105 = 15.5 meter.
Dude
6 years ago
The body first go up with a velocity 4m/s.            The time elapsed in going up:
v=u-gt _______(g is in opposite direction)     =>  t=4/9.8 =0.405 sec.                                     Total time of flight = 0.816 sec.                         Now the body is at the height from which it was dropped but now time left =4.18 sec.          now;  s = ut + 0.5 * g * t2                                   =>  s = 16.73 + 85.61
=>  s = 102.34 ~ 102.5                                     Therefore the body travels 102.5m downward, so the height of the body from the ground = 120.5 – 102.5                           Hence, height of the body from the ground is 18 m.
rohini
4 years ago
s=ut+0.5at2
s= -4×5 + 0.5× 9.8×5×5
.   = 102.5 this is the height above the ground 
so. now 120.5 -102.5=18m
tthis is the height from the ground